OpenStudy (anonymous):

Solve the system by the substitution method. x^2 + y^2 = 113 x + y = 15 A. {( -8, 7), ( -7, 8)} B. {( 8, -7), ( 7, -8)} C. {( 8, 7), ( 7, 8)} D. {( -8, -7), ( -7, -8)}

OpenStudy (nubeer):

x+y = 15 x = 15- y now put this value of x in first equation x^2+y^2 = 113 putting the new value of x (15-y)^2 +y^2 = 113 now just solve it and u will get 2 values of y.

OpenStudy (anonymous):

would it be C?

OpenStudy (nubeer):

i don't know .. i haven't solved.. did you solve this?

OpenStudy (nubeer):

well just did and yes its C

OpenStudy (anonymous):

Thank you very much!

OpenStudy (nubeer):

you are most welcome..